Richard​ Dulski's firm is about to bid on a new radar system. Although the product uses new​ technology, Dulski believes that a learning rate of 90​% is appropriate. The first unit is expected to take 780 ​hours, and the contract is for 14 units.

​a) The total amount of hours to build the 14 units​ = ______ hours​ (round your response to the nearest whole number).

b) The average time to build each of the 14 units = ____
Answer #1

Answer:

b -0.152003093 Unit Time 1 780.00 702.00 N 3 660.04 4 5 631.80 610.73 594.04 6 7 580.28 8 568.62 558.53 9 10 549.66 11 541.75

b = Learning Constant = LN(Learning Percentage)/ LN(2) = LN(0.90)/LN(2) = -0.152

Tn = T1*n^b

Where T1 = 780

Total time to build the 14 units = Sum of time to build unit 1 to 14 = 8362.50 ~ 8363 hours

Average time to build the each of the 14 units = Total time/14 = 597.32 hours